OpenStudy (anonymous):

1. The formula for the amount of simple interest earned is given by the formula I = PRT where I = Interest, P = principal, R = rate, and T = time. Solve this formula for T. Show all steps in your work.

OpenStudy (nurali):

I = PRT Dividing by PR I/PR=PRT/PR T=I/PR

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (anonymous):

Thanks

OpenStudy (nurali):

My Pleasure.